Prove that, for A 1 and A2 disjoint. P(AI UA2 B)-PA B)-P(A2B)
Answer #1

as we know that P(X|Y) =P(X n Y)/P(Y)

also for disjoint events P(A1 u A2)=P(A1)+P(A2)

therefore P(A1 u A2|B)=P((A1 u A2) n B)/P(B)=(P(A1)+P(A2)) n P(B))/P(B)

=(P(A1 n B)/P(B)+(P(A2 n B)/P(B)

P(A1 u A2|B)=P(A1|B)+P(A2|B)

(hence proved)